OpenStudy (anonymous):

please help me big quiz tomorrow!! the arch of the gateshead bridge millenium bridge forms a parabola with equation y=-.016(x-52.5)^2+45 where the x is the horizontal distance in metera from the archs left end and y is the distance in meters from the base of the arch. What is the width of the arch?

OpenStudy (anonymous):

solve for x using the formula \[x=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a }\]

OpenStudy (anonymous):

dont i have to simplify the equation first? or how would i plug it in?

OpenStudy (anonymous):

yes simplify it first

OpenStudy (anonymous):

and after i put it in the equation what do i do?

OpenStudy (anonymous):

how would we find the width?

OpenStudy (anonymous):

y=-.016(x-52.5)^2+45 -.016(x-52.5)^2+45=0 divide by -0.016 to get (x-52.5)^2 - 2812.5=0 x^2 -105x +2756.25-2812.5=0 x^2 -105x +2756.25-2812.5=0 x^2 -105x -56.25=0

OpenStudy (anonymous):

ooh i see...

OpenStudy (anonymous):

x=105.53, x=0.533 width = 105.53-0.533=104.997 = or 105 m

OpenStudy (anonymous):

oooh i think i get it!!!

OpenStudy (anonymous):

ok good correction: sorry the other x=-0.533 its negative

OpenStudy (anonymous):

yea i figured;) thank you very much!!

OpenStudy (anonymous):

ok yw...good luck now :D

OpenStudy (anonymous):

thank you!!!